OpenStudy (anonymous):

how do i solve this problem? Find the general equation of the line; that is, write the equation in the form Ax+By=C Slope undefined; containing the point (4,3)

OpenStudy (amistre64):

what does an undefinned slope look like? mathically that is

OpenStudy (amistre64):

would you agree that n/0 is undefined for any n?

OpenStudy (amistre64):

if so, then use it in the point slope format and algebra it out into the general format: y - yo = n/0 (x-xo) 0(y - yo) = n (x-xo) 0y - 0 = nx -n xo -nx + 0y = -n xo what value of n will satisfy?

OpenStudy (anonymous):

i honestly have no idea what it looks like